The Donald E. Knuth Scholarship: 1994 Scholar and 1995 Announcement

At the 15th Annual Meeting of TUG, Shelley-Lee Ames was honored as the 1994 Donald E. Knuth Scholarship winner. Shelley works at the University of Manitoba for the Canadian Mathematical Society (Société mathématique du Canada) where she prepares, formats and proofs all papers published by the society in their Journal and Bulletin.

Announcement of the 1995 competition

One Knuth Scholarship will be available for award in 1995. The competition will be open to all TEX users holding support positions that are secretarial, clerical or editorial in nature. It is therefore not intended for those with a substantial training in technical, scientific or mathematical subjects and, in particular, it is not open to anyone holding, or studying for, a degree with a major or concentration in these areas.

The award will consist of an expense-paid trip to the 1995 TUG Annual Meeting at St. Petersburg, Florida, and to the Scholar’s choice from the short courses offered in conjunction with that meeting; and TUG membership for 1995, if the Scholar is not a TUG member, or for 1996, if the Scholar is already a TUG member. A cap of $2000 has been set for the award; however, this does not include the meeting or course registration fee, which will be waived.

To enter the competition, applicants should submit to the Scholarship Committee, by the deadline specified below, the input file and final TEX output of a project that displays originality, knowledge of TEX, and good TEXnique.

The project as submitted should be compact in size. If it involves a large document or a large number of documents then only a representative part should be submitted, together with a description of the whole project. For example, from a book just one or two chapters would be appropriate.

The project may make use of a macro package, either a public one such as L^{A}TEX or one that has been developed locally; such a macro package should be identified clearly. Such features as sophisticated use of math mode, of macros that require more than “filling in the blanks”, or creation and use of new macros will be taken as illustrations of the applicant’s knowledge.

All macros created by the candidate should be well documented with clear descriptions of how they should be used and an indication of how they work internally.

All associated style files, macro-package files, etc., should be supplied, or a clear indication given of any widely available ones used (including version numbers, dates, etc.); clear information should be provided concerning the version of TEX used and about any other software (e.g. particular printer drivers) required. Any nonstandard fonts should be identified and provided in the form of .tfm and .pk files suitable for use on a 300dpi laser printer.

While the quality of the typographic design will not be an important criterion of the judges, candidates are advised to ensure that their printed output adheres to sound typographic standards; the reasons for any unusual typographic features should be clearly explained.

All files and documents comprising the project must be submitted on paper; the input files should be provided in electronic form as well. Suitable electronic media are IBM PC-compatible or Macintosh diskettes, or a file sent by electronic mail.

A brochure with additional information is available from the TUG office. To obtain a copy, or to request instructions on e-mail submission, write to the address at the end of this announcement, or send a message by e-mail to tug@tug.org with the subject “Knuth Scholarship request”.

Along with the project, each applicant should submit a letter stating the following:

1. affirmation that he/she will be available to attend the 1995 TUG Annual Meeting;

2. affirmation of willingness to participate on the committee to select the next Scholar. Each applicant should also submit a curriculum vitae summarizing relevant personal information, including:

1. statement of job title, with a brief description of duties and responsibilities;

2. description of general post-secondary school education, TEX education, identifying courses attended, manuals studied, personal instruction from experienced TEX users, etc.;

3. description of TEX resources and support used by the candidate in the preparation of the project.

Neither the project nor the curriculum vitae should contain the applicant’s name or identify the applicant. These materials will be reviewed by the committee without knowledge of applicants’ identities. If, despite these precautions, a candidate is identifiable to any judge, then that judge will be required to make this fact known to the others and to the TUG board members responsible for the conduct of the judging.

The covering letter, curriculum vitae, and all macro documentation that is part of the project input should be in English. (English is not required for the output of the project.) However, if English is not the applicant’s native language, that will not influence the decision of the committee.

Selection of the Scholarship recipient will be based on the project submitted.

Schedule

The following schedule will apply (all dates are in 1995):

7 April Deadline for receipt of

submissions

21 April–2 June Judging period

9 June Notification of winner 24–28 July 1995 Annual Meeting,

St. Petersburg, Florida

The 1995 Scholarship Committee consists of • Chris Rowley, Open University, UK (Chair);

• David Salomon, California State University, Northridge, USA;

• Shelly-Lee Ames, University of Manitoba, Canada.

Where to write

All applications should be submitted to the Committee in care of the TUG office:

TEX Users Group

Attn: Knuth Scholarship Competition

P. O. Box 869

Santa Barbara, CA 93102 USA

    e-mail: tug@tug.org

Nico Poppelier

Liaison to the Donald E. Knuth Scholarship Committee